OpenStudy (anonymous):

If K = {(x, y ) | x - y = 5}, is Set K a function? Yes No

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

I have absolutely no idea.

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrcircle.htm this help me out when i was doing these types of questions

OpenStudy (freckles):

Basically it asking you if the linear equation x-y=5 passes the vertical line test

OpenStudy (anonymous):

I don't know! Lol

OpenStudy (freckles):

do you know what x-y=5 looks like?

OpenStudy (anonymous):

No

OpenStudy (freckles):

do you know y=mx+b is a linear equation with y-intercept b and slope m?

OpenStudy (anonymous):

I haven't done slope in 2 years

OpenStudy (freckles):

x-y=5 -y=-x+5 (subtracted x on both sides) y=x-5 (multiplied -1 on both sides) y=x-5 is a line with slope 1 and y intercept -5

OpenStudy (freckles):

can you draw this now?

OpenStudy (freckles):

 this is what that line looks like does it pass the vertical line test?